OpenStudy (anonymous):

Which is the equation of the line, in point slope form, that has a slope of 2 and passes through the point (-8, 1)?

OpenStudy (anonymous):

y=m*x + b 1=2*(-8) + b Solve for b. m was given in the problem statement.